OpenStudy (anonymous):

for a system of equations, what kind of solution would 12=-1 be????????? PLEASE HELPPPPPPP

jimthompson5910 (jim_thompson5910):

12=-1 is false because 12 and -1 are two completely different numbers. So assuming you followed the steps properly, this means that there are no solutions to this system. This system is called "inconsistent".

OpenStudy (anonymous):

Thanks man

jimthompson5910 (jim_thompson5910):

Example: y = x+2 y = x+3 ----------- y = x+2 x+3 = x+2 ... plug in y = x+3 x-x = 2-3 0 = -1 ... false, so there are no solutions If you graphed y=x+2 and y=x+3, you would find that the two lines never intersect. They are parallel lines. You need a point of intersection to have a solution.

jimthompson5910 (jim_thompson5910):

no problem